# Learning from Neighbors with PHIBP: Predicting Infectious Disease Dynamics in Data-Sparse Environments

### v1

@@ 2.1 Predicting the unseen and alpha/beta diversities
We now recount some of the novel developments in [5] that will figure prominently in our analysis. We will use this compound Poisson form, together with the posterior local rates $\tilde{\sigma}_{j,l}(H_{l})$, to define and compute our Bayesian alpha- and beta-diversity measures and disease prediction functionals. Specifically the representation in (2.3) indicates there is a decomposition of the pairs of rates and species labels $(\lambda_{l},Y_{l})_{l\geq 1}$ in terms of unobserved species and corresponding rates $(\lambda^{\prime}_{l},Y^{\prime}_{l})_{l\geq 1}$ and the $\varphi$ observed species $\tilde{Y}_{\ell}$ and corresponding rates $H_{\ell}$ for $\ell\in[\varphi].$ The localized posterior rates $\tilde{\sigma}_{j,l}(H_{l})$ allow us to construct novel Bayesian analogues of classical $\alpha$- and $\beta$-diversity metrics that account for uncertainty and unobserved species.
Within this framework, among other possibilities, we define alpha-diversity (within-group diversity) via the Shannon entropy, which is the random variable:
@@
For beta-diversity (between-group diversity), we define a dissimilarity based on the Bray-Curtis index, which is the random variable:
@@
Our constructs while new are in the form of more classical measures of diversity as detailed in [9, 10]. In ecological and epidemiological settings, alpha-diversity measures the within-region richness and evenness of types—in our case, how many diseases circulate in a county and how evenly their burdens are distributed—while beta-diversity quantifies between-region dissimilarity, capturing how disease profiles differ across counties or how transmission patterns diverge geographically. Classically, these metrics are computed from observed proportions, but such compositional methods break down in the presence of many zeros, causing artificial inflation of dissimilarity when low-prevalence types are simply unobserved. The PHIBP framework replaces fixed proportions with latent abundance rates, producing alpha- and beta-diversities that properly distinguish sparse detection from true absence. Moreover, these quantities are Bayesian random variables, yielding full posterior distributions that reflect uncertainty—critical when data are sparse or uneven across regions—rather than single point estimates. This same decomposition underlies PHIBP’s approach to the unseen-disease problem, enabling principled prediction of diseases not yet observed but likely to appear under continued sampling. In our previous PHIBP experiments on the Dorado Outcrop microbiome data, which is based on data originally investigated by [7, 10], this Bayesian, rate-based treatment proved essential: the model captured rare shared taxa, avoided pseudocount artifacts, and produced more stable alpha- and beta-diversity estimates than compositional alternatives. The same advantages carry over directly to the disease-prediction setting studied here.
@@ 4.2 Practical Extensions for Disease Surveillance
Alternative disease spaces: The beta-diversity patterns we observed arose from geographic proximity, but the framework can incorporate other distance metrics—genomic similarity between strains, human mobility networks, or environmental factors—by choosing appropriate Lévy measures $\tau_{j}$ that encode domain-specific transmission dynamics [4, Section 7.5.3].
@@ 4.4 Implications for Real-Time Implementation
The same architecture that allows borrowing of statistical strength across California’s 58 counties could be applied to multi-site clinical trials, environmental monitoring networks, or financial contagion modeling—each instantiating the same fundamental duality with domain-appropriate specifications [4, Section 7.5].

### v2

@@ 3.1 Predicting the unseen and alpha/beta diversities
We now recount some of the novel developments in [6] that will figure prominently in our analysis. The representation in (3.3) decomposes the pairs of rates and species labels $(\lambda_{l},Y_{l})_{l\geq 1}$ into two components: unobserved species with rates $(\lambda^{\prime}_{l},Y^{\prime}_{l})_{l\geq 1}$ and the $\varphi$ observed species $\tilde{Y}_{\ell}$ with posterior mean rates $H_{\ell}$ for $\ell\in[\varphi]$. We use this compound Poisson representation, together with the posterior local rates $\sigma_{j,\ell}(H_{\ell})$, to define and compute Bayesian alpha- and beta-diversity measures and disease prediction functionals.
Within this framework, we define alpha-diversity (within-group diversity) via the Shannon entropy:
@@
For beta-diversity (between-group diversity), we define a dissimilarity based on the Bray-Curtis index:
@@
Our constructs, while new, are in the form of more classical measures of diversity as detailed in [11, 14]. In ecological and epidemiological settings, alpha-diversity measures the within-region richness and evenness of types—in our case, how many diseases circulate in a county and how evenly their burdens are distributed—while beta-diversity quantifies between-region dissimilarity, capturing how disease profiles differ across counties or how transmission patterns diverge geographically.
Classically, these metrics are computed from observed proportions, but such compositional methods break down in the presence of many zeros, causing artificial inflation of dissimilarity when low-prevalence types are simply unobserved. The PHIBP framework replaces fixed proportions with latent abundance rates, producing alpha- and beta-diversities that properly distinguish sparse detection from true absence. Moreover, these quantities are Bayesian random variables, yielding full posterior distributions that reflect uncertainty—critical when data are sparse or uneven across regions—rather than single point estimates.
This same decomposition underlies PHIBP’s approach to the unseen-disease problem, enabling principled prediction of diseases not yet observed but likely to appear under continued sampling. The random variable $\varphi$ in (3.3) represents the total number of disease types that could appear in the observed data, including both those actually detected and those present in the environment but unobserved due to finite sampling. The posterior distribution of the unobserved rates $(\lambda^{\prime}_{l})_{l\geq 1}$ provides a basis for predicting the prevalence and regional distribution of diseases that have not yet been observed but may emerge with additional surveillance effort or in future samples.
In our previous PHIBP experiments on the Dorado Outcrop microbiome data, originally investigated by [8, 14], this Bayesian, rate-based treatment proved essential: the model captured rare shared taxa, avoided pseudocount artifacts, and produced more stable alpha- and beta-diversity estimates than compositional alternatives. The same advantages carry over directly to the disease-prediction setting studied here.
@@ 5.2 Practical Extensions for Disease Surveillance
Alternative disease spaces: The beta-diversity patterns we observed arose from geographic proximity, but the framework can incorporate other distance metrics—genomic similarity between strains, human mobility networks, or environmental factors—by choosing appropriate Lévy measures $\tau_{j}$ that encode domain-specific transmission dynamics [5, Section 7.5.3].
@@ 5.4 Implications for Real-Time Implementation
The same architecture that allows borrowing of statistical strength across California’s 58 counties could be applied to multi-site clinical trials, environmental monitoring networks, or financial contagion modeling—each instantiating the same fundamental duality with domain-appropriate specifications [5, Section 7.5].